following statement is true and which are false? Give reasons for your answers, in the form of a short proof or a counter example.





Every continuous function is differentiable

1
Expert's answer
2021-12-27T03:27:34-0500

False.

For example, the function "f(x)=|x|" is continuous at "x=0" :


"\lim\limits_{x\to 0^-}f(x)=0=\lim\limits_{x\to 0^+}f(x)=>\lim\limits_{x\to 0}f(x)=0"

"f(0)=|0|=0=\lim\limits_{x\to 0}f(x)"

On the other hand,


"\lim\limits_{h\to 0^-}\dfrac{f(0+h)-f(0)}{h}=\lim\limits_{h\to 0^-}\dfrac{|h|-0}{h}=-1"

"\lim\limits_{h\to 0^+}\dfrac{f(0+h)-f(0)}{h}=\lim\limits_{h\to 0^+}\dfrac{|h|-0}{h}=1"

"\lim\limits_{h\to 0^-}\dfrac{f(0+h)-f(0)}{h}=-1"

"\not=1=\lim\limits_{h\to 0^+}\dfrac{f(0+h)-f(0)}{h}"

Then "\lim\limits_{h\to 0}\dfrac{f(0+h)-f(0)}{h}" does not exist.

Therefore the function "f(x)=|x|" is not differentiable at "x=0."